OpenStudy (mallorysipp234):

f(x+h)-f(x)/h f(x)=2x^2-5x how does this H play in??

OpenStudy (mallorysipp234):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

let's find what f(x+h) is equal to start with f(x)=2x^2-5x then replace every copy of x with (x+h) so we'll then have f(x+h)=2(x+h)^2-5(x+h) making sense so far?

OpenStudy (mallorysipp234):

wait. so f(x+h)-2(x+h)^2-5(x+h)/h, ok yes that makes sense.

jimthompson5910 (jim_thompson5910):

we're focusing on just f(x+h) for now. Don't worry about [f(x+h)-f(x)]/h

jimthompson5910 (jim_thompson5910):

Do you see how I went from f(x)=2x^2-5x to f(x+h)=2(x+h)^2-5(x+h) ??

OpenStudy (mallorysipp234):

ohh, you even replaced the f(x) with x+h, too. Yes i do see where you're coming from

jimthompson5910 (jim_thompson5910):

yes every x gets replaced with x+h even the first x so we go from f(x) to f(x+h)

OpenStudy (mallorysipp234):

yes yes

jimthompson5910 (jim_thompson5910):

now let's simplify the right side of f(x+h)=2(x+h)^2-5(x+h)

jimthompson5910 (jim_thompson5910):

how would you simplify 2(x+h)^2-5(x+h) ?

jimthompson5910 (jim_thompson5910):

or I guess I should say "expand out"

OpenStudy (mallorysipp234):

oh ok, so we're multiplying 2 by (x+h)^2 and -5 by (x+h)

jimthompson5910 (jim_thompson5910):

what happens when you FOIL out (x+h)^2 ?

OpenStudy (mallorysipp234):

x^2+xh+xh+h^2

jimthompson5910 (jim_thompson5910):

which becomes x^2+2xh+h^2

OpenStudy (mallorysipp234):

okay now multiply all those terms by 2?

jimthompson5910 (jim_thompson5910):

\[\Large f(x+h) = 2(x+h)^2-5(x+h)\] \[\Large f(x+h) = 2(x^2+2xh+h^2)-5(x+h)\] now distribute

OpenStudy (mallorysipp234):